# TBRD: TESLA Authenticated UAS Broadcast Remote ID

### v1

@@ Abstract
Mysterious sightings of Unmanned Aircraft Systems (UAS) over U.S. military facilities, suburban neighborhoods, and commercial airports have intensified scrutiny of drone activity. To increase accountability, the Federal Aviation Administration (FAA) introduced a Remote ID mandate, requiring unmanned aircraft to broadcast their location, operator’s location, and identity in real-time. However, current standards leave authentication mechanisms underspecified, enabling spoofing, relay, and replay attacks that can undermine surveillance efforts and potentially disrupt UAS-to-UAS coordination in future deployments. In this paper, we propose TBRD, a practical system for authenticating Remote ID messages in a manner that aligns with existing standards and UAS capabilities. TBRD leverages the TESLA protocol and mobile device TEEs, and introduces a verification mechanism to build a lightweight, mission-scoped authentication system that is both computationally efficient and requires a low communication footprint. We evaluate the performance of TBRD using both an FAA-requirements compatible proof-of-concept implementation for performance metrics and a simulated 4-drone swarm mission scenario to demonstrate its security guarantees under adversarial conditions. Our system provides a 50% reduction in authentication overhead compared to digital signatures and a 100x reduction in computation time. Our results demonstrate that TBRD can be integrated into current Remote ID infrastructures to provide a scalable, standards-compliant message authentication for both regulatory and operational use cases.
I Introduction
@@
In recent years, the rapid proliferation of Unmanned Aircraft Systems (UASs) has raised pressing concerns around safety, accountability, and regulatory enforcement. Reports of unidentified drones operating near critical infrastructure—such as military bases, commercial airports, and densely populated urban areas—have increased significantly, heightening fears of both accidental collisions and deliberate misuse [1, 2, 3, 4]. Law enforcement and airspace regulators face difficulty identifying and holding remote pilots accountable, especially when drone operations lack reliable mechanisms for real-time identification. In response, the U.S. Federal Aviation Administration (FAA) introduced legislation requiring drones to broadcast their location, operator information, and unique identifier in real time, a mandate commonly referred to as the Remote ID rule [5].
@@
Despite its regulatory importance, the Remote ID standard remains critically incomplete from a security perspective. Most notably, Remote ID broadcasts are unauthenticated, meaning any observer has no way to determine whether a message truly originated from a legitimate UAS. This opens the door to spoofing, relay, and replay attacks that can be executed using simple off-the-shelf tools [6, 7, 8]. For example, an adversary could replay previously captured broadcasts near a restricted area, triggering false alarms or framing legitimate operators. Such vulnerabilities undermine the integrity of Remote ID data, erode trust in enforcement systems, and, in the future, may even interfere with UAS-to-UAS coordination in decentralized swarm operations.
@@
While digital signatures offer a natural approach to authenticating broadcasts, they introduce practical challenges in UAS deployments. Public key cryptography produces large signature outputs, adding nontrivial overhead to Remote ID messages that are already size-constrained by broadcast standards. Frequent signing also imposes computational burdens on resource-limited UAS hardware. Large-scale key management, revocation, and distribution further complicate deployment, even with assistance from a backend service. In contrast, the Timed-Efficient Stream Loss-Tolerant Authentication (TESLA) [9] protocol provides similar security guarantees using symmetric cryptography and delayed key disclosure, significantly reducing both computational and communication costs. These properties make TESLA especially suitable for intermittent, real-time systems like Remote ID.
In this paper, we propose TBRD, a system for authenticated Remote ID broadcasting that is compatible with existing standards and realistic deployment constraints. TBRD leverages TESLA for efficient broadcast authentication, uses Trusted Execution Environments (TEEs) on mobile devices to securely manage keychains, and introduces a UAS Service Supplier (USS) component to manage key commitments and enable delayed key validation. The use of TESLA also removes the need for continuous connectivity to a backend: observers can passively collect broadcasts and verify them later once keys are disclosed. TEEs are increasingly present in commercial mobile phones and even some drone controllers. In cases where they are unavailable, our architecture gracefully falls back to secure local key storage. This flexibility enables TBRD to scale with the evolving drone ecosystem while maintaining low overhead and high verifiability.
@@
While TESLA has been adopted in systems like Galileo GNSS [10], applying it to Remote ID requires a fundamentally different design. GNSS satellites are few in number, their identities and codes are fixed, and their public keys rarely change; new satellites are added only every few years. In contrast, drones are purchased, deployed, and retired dynamically, with new Remote ID transmitters appearing unpredictably. Remote ID authentication must therefore accommodate an open and evolving population of devices while ensuring that observers can validate messages from unknown transmitters without prior provisioning. TBRD addresses this challenge by enabling verifiable broadcasts from previously unknown UASs. It utilizes delayed key disclosure and public commitments stored at the USS, allowing the observers to authenticate messages offline once the corresponding keys become available.
@@
Specifically, we make the following contributions. We design TBRD, a Remote ID authentication system that utilizes the TESLA broadcast authentication protocol within the drone ecosystem, enabling verifiable, spoof-resistant broadcasts with minimal computational and communication overhead. We show how mission-scoped keychains can be securely generated and managed using TEEs on commodity mobile devices, reducing the risk of key compromise and enabling deployment on resource-constrained UAS platforms. We integrate a UAS Service Supplier (USS) into the authentication workflow to distribute key commitments and mission metadata, allowing observers to authenticate messages using TESLA-disclosed keys without prior provisioning or persistent connectivity. Our implementation, built on OpenDroneID and Android TEEs, adheres to FAA standards. We compare TBRD against digital signature-based approaches in terms of signing cost, message size, and energy consumption. Finally, using a custom high-fidelity 4-UAS swarm simulation with a decentralized collision avoidance algorithm, we show how TBRD mitigates spoofing, replay, and relay attacks that can otherwise disrupt coordinated flight. Our system provides a 50 percent reduction in authentication overhead compared to digital signatures, requires a fraction of the computational effort — approximately a (or 100×) reduction in signing time from 1.2 seconds to 10 milliseconds. By enabling scalable and lightweight authentication within existing broadcast constraints, we are actively involved in proposing the integration of TBRD into the broader Drone Remote Identification Protocol (DRIP) framework  [11].
II Background
@@
The FAA and Remote ID: The Federal Aviation Administration (FAA) is responsible for overseeing UAS operations in the United States. Historically, rules have been drafted to ensure public safety in an increasingly congested National Airspace System (NAS). These efforts have been heavily influenced by both UAS manufacturers and hobbyists alike. New policies have at times led to friction and significant public scrutiny [12]. Nonetheless, 14 CFR Part 107 establishes rules for UAS operations, remote pilot certification requirements, and rules for flying UAS over human beings [13]. To hold remote pilots accountable, the FAA now enforces 14 CFR Part 89, which requires operators to comply with the rule on Remote ID.
@@
Pilots can comply with Remote ID in 3 different ways. First, an operator can operate a UAS that has been equipped by its manufacturer with the capability to broadcast identification and location information. Second, an operator can modify their existing UAS to add the ability to broadcast Remote ID information. Finally, an operator can fly within an FAA-recognized identification area (FRIA) without the requirement to broadcast any Remote ID information. In either of the first two scenarios, the UAS ID, the operator’s position, and the UAS’s position and velocity information must be broadcast [14].
@@
The regulations on Remote ID do not dictate the system design of a Remote ID transponder, nor do they require any means of message authentication. It only specifies the functional and performance requirements that any system must meet. In other words, the actual design and development of a Remote ID transponder is left to UAS manufacturers and, in the case of open-source solutions, UAS operators. The UAS manufacturer DJI produces systems that are Remote ID compliant through the use of proprietary protocols called OcuSync. This capability has been included even before the rule on Remote ID was codified into US law. While it remains closed-source, legacy versions of their system, called DroneID, have been reverse-engineered to uncover significant security flaws [15]. Due to the complexities of reverse engineering proprietary implementations, our work aims to enhance open source solutions for Remote ID compliance.
@@
One such effort to promote an open source solution for Remote ID is the publication of the Standard Specification for Remote ID and Tracking, F3411-22a, by the American Society for Testing and Materials (ASTM) [16]. This specification provides the technical details that standardize manufacturer approaches and open source efforts to build Remote ID-compliant systems. The main open source effort that complies with ASTM F3411-22a is the project OpenDroneID [17]. In addition to offering an open-source C library that individual developers can use to build Remote ID systems, other open-source UAS ground station control developers, such as ArduPilot and MAVLink, offer tools and software to support OpenDroneID integration. This has enabled several manufacturers to build and offer relatively inexpensive add-on modules that can retrofit a UAS with Remote ID capabilities. Despite the lack of regulations requiring any level of security for the authentication of Remote ID, the ASTM standard outlines a mechanism for sending authentication data. The specifics of the ASTM message structure will be discussed in greater depth in section III.
@@
The TESLA protocol: Broadcast authentication has proved to be a complex issue to solve for several years. Many public services, such as the Global Positioning Service (GPS), are broadcast via some wireless medium. These services are vulnerable to growing threats of jamming, spoofing, or record-and-replay attacks. As a result, many works have attempted to provide an efficient means for receivers to verify and authenticate broadcast transmissions. One such method is TESLA [9]. The motivating concept of the TESLA protocol is that each broadcast packet includes a Message Authentication Code (MAC) computed with a key known only to the transmitter. That key is disclosed at a later time interval, whereby receivers can authenticate previously broadcast messages. TESLA authentication is currently in use by the Galileo global navigation satellite system. The Galileo Open Service Navigation Message Authentication (OSNMA) officially became operational in July of 2025 and utilizes a combination of the TESLA protocol and digital signature to authenticate satellite positioning messages [18, 10]. Furthermore, the TESLA authentication has been proposed as a means to authenticate various types of broadcast services, including Automated Identification System (AIS) for marine traffic information broadcast [19], as well as Automated Dependent Surveillance Broadcast (ADS-B) systems for air traffic information broadcast [20]. More recently, it was identified as a potential solution for providing authentication for the very-high-frequency data exchange system (VDES) that is anticipated to enhance a wide variety of maritime communications capabilities  [21]. We are the first to propose using TESLA for Remote ID broadcast authentication.
@@
TEE capabilities: TEEs in mobile computing are secure, isolated environments where the most critical operations related to key generation and storage can be safely executed [22]. Many mobile applications are becoming increasingly reliant on the services provided by TEEs. They are used to provide services ranging from Mobile Network Operator (MNO) billing, digital rights management, and biometric authentication [23, 24]. In our specific application, the TEE’s integrity and secure storage capabilities are leveraged to conduct the transmission setup procedures required by the TESLA protocol.
III Design of TBRD
@@
III-A Threat Model and Assumptions
Our threat model addresses adversaries who have radio access to Remote ID broadcasts and aim to exploit the absence of authentication to inject false or misleading Remote ID broadcasts. We assume an attacker can either passively observe and/or actively transmit wireless data with basic SDR capabilities (e.g., USRP or HackRF). Furthermore, they have sufficient knowledge to record and replay Wi-Fi beacons or Bluetooth advertisements, the two FAA-approved broadcast media. We also consider attackers who can craft and transmit arbitrary Remote ID messages using open-source toolchains, including those that comply with the Remote ID standard but contain spoofed data. Figure 1 depicts an example scenario where the attacker captures Remote ID broadcasts from a legitimate UAS mission and replays them at a later time to create the appearance of unauthorized activity. For example, during a public event like the Super Bowl in New Orleans, temporary flight restrictions may be in place. The attacker replays previously captured Remote ID messages in the restricted zone, making it appear as though a registered operator is conducting an illegal flight. This can trigger enforcement actions or initiate investigations against legitimate operators. In coordinated attacks, repeated replays across multiple zones could be used to overwhelm response efforts and misdirect law enforcement attention.
Figure 1: Manipulation of Broadcast Remote ID. An attacker intercepts legitimate Remote ID broadcasts from a compliant UAV, modifies the captured identification and telemetry data, and retransmits the spoofed information to deceive observers.
@@ III-A Threat Model and Assumptions
III-B System Overview

### v2

# TBRD: Tesla Authenticated Broadcast Remote ID

@@ Abstract.
Mysterious sightings of Unmanned Aircraft Systems (UAS) over U.S. military facilities, suburban neighborhoods, and commercial airports have intensified scrutiny of drone activity. To increase accountability, the Federal Aviation Administration (FAA) introduced a Remote ID mandate, which requires unmanned aircraft to broadcast their location, the location of the operator, and their identity in real-time. However, current standards leave authentication mechanisms unspecified, enabling spoofing, relay, and replay attacks that can undermine surveillance efforts and potentially disrupt UAS-to-UAS coordination in future deployments. In this paper, we propose TBRD, a practical system for authenticating Remote ID messages in a manner that aligns with existing standards and UAS capabilities. TBRD leverages the TESLA protocol and mobile device TEEs, and introduces a verification mechanism to build a lightweight, mission-scoped authentication system that is both computationally efficient and requires a low communication footprint. We evaluate the performance of TBRD using both an FAA-requirements compatible proof-of-concept implementation for performance metrics and a simulated 4-drone swarm mission scenario to demonstrate its security guarantees under adversarial conditions. Our system provides a 50% reduction in authentication overhead compared to digital signatures and a 100x reduction in computation time. Our results demonstrate that TBRD can be integrated into current Remote ID infrastructures to provide a scalable, standards-compliant message authentication for both regulatory and operational use cases.
copyright: noneconference: ; ;
1. Introduction
@@
In recent years, the rapid proliferation of Unmanned Aircraft Systems (UASs) has raised pressing concerns around safety, accountability, and regulatory enforcement. Reports of unidentified drones operating near critical infrastructure—such as military bases, commercial airports, and densely populated urban areas—have increased significantly, heightening fears of both accidental collisions and deliberate misuse (muntean_faa_2025; bonavita_foreign_2025; khalil_mysterious_2024; whitaker_drone_2025). Law enforcement and airspace regulators face difficulty identifying and holding remote pilots accountable, especially when drone operations lack reliable mechanisms for real-time identification. In response, the U.S. Federal Aviation Administration (FAA) introduced legislation requiring drones to broadcast their location, operator information, and unique identifier in real time, a mandate commonly referred to as the Remote ID rule (federal_aviation_administration_faa_2019).
@@
Despite its regulatory importance, the Remote ID standard remains critically incomplete from a security perspective. Most notably, Remote ID broadcasts are unauthenticated, meaning any observer has no way to determine whether a message truly originated from a legitimate UAS. Both Remote ID and Automated Dependent Surveillance-Broadcast (ADS-B) exemplify a fundamental trade-off in aviation safety systems where the priority on open, accessible broadcasts for collision avoidance and airspace awareness inherently creates security vulnerabilities (khan_survey_2025). Adding authentication or encryption increases complexity, can reduce interoperability, and potentially compromise the real-time monitoring capabilities that make these systems effective for their primary safety mission. This opens the door to spoofing, relay, and replay attacks that can be executed using simple off-the-shelf tools (jet_jjshootsremoteidspoofer_2025; tsuboi_drone_2024; muller_cyber-defence-campusdroneremoteidspoofer_2025). For example, an adversary could replay previously captured broadcasts near a restricted area, triggering false alarms or framing legitimate operators. Such vulnerabilities undermine the integrity of Remote ID data, erode trust in enforcement systems, and, in the future, may even interfere with UAS-to-UAS coordination in decentralized swarm operations.
@@
While digital signatures offer a natural approach to authenticating broadcasts, they introduce practical challenges in UAS deployments. Public key cryptography produces large signature outputs, adding nontrivial overhead to Remote ID messages that are already size-constrained by broadcast standards. Frequent signing also imposes computational burdens on resource-limited UAS hardware. Large-scale key management, revocation, and distribution further complicate deployment, even with assistance from a backend service. In contrast, the Timed-Efficient Stream Loss-Tolerant Authentication (TESLA) (perrig_tesla_2002) protocol provides similar security guarantees using symmetric cryptography and delayed key disclosure, significantly reducing both computational and communication costs. These properties make TESLA especially suitable for intermittent, real-time systems like Remote ID.
@@
TESLA has been adopted in systems like Galileo GNSS (european_gnss_supervisory_authority_galileo_2022). However, applying it to Remote ID requires a fundamentally different design. GNSS satellites are few in number, their identities and codes are fixed, and their public keys rarely change; new satellites are added only every few years. In contrast, drones are purchased, deployed, and retired dynamically, with new Remote ID transmitters appearing unpredictably. Remote ID authentication must therefore accommodate an open and evolving population of devices while ensuring that observers can validate messages from unknown transmitters without prior provisioning. TBRD addresses this challenge by enabling verifiable broadcasts from previously unknown UASs. It utilizes delayed key disclosure and public commitments stored at the USS (UAS Service Supplier), allowing the observers to authenticate messages offline once the corresponding keys become available.
In this paper, we propose TBRD, a system for authenticated Remote ID broadcasting that is compatible with existing standards and realistic deployment constraints. TBRD leverages TESLA for efficient broadcast authentication, uses Trusted Execution Environments (TEEs) on mobile devices to securely manage keychains, and introduces a UAS Service Supplier (USS) component to manage key commitments and enable delayed key validation. The use of TESLA also removes the need for continuous connectivity to a backend: observers can passively collect broadcasts and verify them later once keys are disclosed. TEEs are increasingly present in commercial mobile phones. In cases where they are unavailable, our architecture gracefully falls back to secure local key storage. This flexibility enables TBRD to scale with the evolving drone ecosystem while maintaining low overhead and high verifiability.
@@
Specifically, we make the following contributions. We design TBRD, a Remote ID authentication system that utilizes the TESLA broadcast authentication protocol within the drone ecosystem, enabling verifiable, spoof-resistant broadcasts with minimal computational and communication overhead. We show how mission-scoped keychains can be securely generated and managed using TEEs on commodity mobile devices, reducing the risk of key compromise and enabling deployment on resource-constrained UAS platforms. We integrate a UAS Service Supplier (USS) into the authentication workflow to distribute key commitments and mission metadata, allowing observers to authenticate messages using TESLA-disclosed keys without prior provisioning or persistent connectivity. Our implementation, built on OpenDroneID and Android TEEs, adheres to FAA standards. We compare TBRD against digital signature-based approaches in terms of signing cost, message size, and energy consumption. Finally, using a custom high-fidelity 4-UAS swarm simulation with a decentralized collision avoidance algorithm, we show how TBRD mitigates spoofing, replay, and relay attacks that can otherwise disrupt coordinated flight. Our system provides a 50 percent reduction in authentication overhead compared to digital signatures and requires a fraction of the computational effort — approximately a 100× reduction in signing time from 1.2 seconds to 10 milliseconds.
2. Background
@@
14 CFR Part 107 establishes rules for UAS operations, remote pilot certification requirements, and rules for flying UASs over human beings (federal_aviation_administration_small_2016). To ensure remote pilots are operating UASs safely, the FAA now enforces 14 CFR Part 89, which requires operators to comply with the rule on Remote ID.
@@
Pilots can comply with Remote ID in 3 different ways. First, a pilot can operate a UAS that has been equipped with the capability to broadcast identification and location information by the manufacturer. Second, an operator can modify their existing UAS to add the ability to broadcast Remote ID information. Finally, an operator flying within an FAA-recognized identification area (FRIA) satisfies the Remote ID requirement without needing to broadcast Remote ID information. In either of the first two scenarios, the UAS ID, the operator’s position, and the UAS’s position and velocity information must be broadcast (federal_aviation_administration_remote_2021).
@@
The regulations on Remote ID do not dictate the system design of a Remote ID transponder, nor do they require any means of message authentication. It only specifies the functional and performance requirements that any system must meet. In other words, the actual design and development of a Remote ID transponder is left to UAS manufacturers and, in the case of open-source solutions, UAS operators. The UAS manufacturer DJI produces systems that are Remote ID compliant through the use of proprietary protocols called OcuSync. This capability has been included even before the rule on Remote ID was codified into US law. While it remains closed-source, legacy versions of their system, called DroneID, have been reverse-engineered to uncover significant security flaws (schiller_drone_2023). Due to the complexities of reverse engineering proprietary implementations, our work aims to enhance open source solutions for Remote ID compliance.
@@
One such effort to promote an open source solution for Remote ID is the publication of the Standard Specification for Remote ID and Tracking, F3411-22a, by the American Society for Testing and Materials (ASTM) (f38_committee_specification_2022). This specification provides the technical details that standardize manufacturer approaches and open source efforts to build Remote ID-compliant systems. The specification outlines an optional digital signature-based authentication mechanism. However, the approach requires persistent internet connectivity for observers to contact a Broadcast Authentication Verifier Service in order to determine message validity. The main open source effort that complies with ASTM F3411-22a is the project OpenDroneID (open_drone_id_open_2025). In addition to offering an open-source C library that individual developers can use to build Remote ID systems, other open-source UAS ground station control developers, such as ArduPilot and MAVLink, offer tools and software to support OpenDroneID integration. This has enabled several manufacturers to build and offer relatively inexpensive add-on modules that can retrofit a UAS with Remote ID capabilities (holybro_remote_2025; blue_mark_dronebeacon_2025). Despite the lack of regulations requiring any level of security or the authentication of Remote ID, the ASTM standard outlines a mechanism for broadcasting authentication messsage data. The specifics of the ASTM message structure will be discussed in greater depth in section 3.
@@
The TESLA protocol: Broadcast authentication has proved to be a complex issue to solve for several years. Many public services, such as the Global Positioning Service (GPS), are broadcast via some wireless medium. These services are vulnerable to growing threats of jamming, spoofing, or record-and-replay attacks. As a result, many works have attempted to provide an efficient means for receivers to verify and authenticate broadcast transmissions. One such method is TESLA (perrig_tesla_2002). The motivating concept of the TESLA protocol is that each broadcast packet includes a Message Authentication Code (MAC) computed with a key known only to the transmitter. That key is disclosed at a later time interval, whereby receivers can authenticate previously broadcast messages. TESLA authentication is currently in use by the Galileo global navigation satellite system. The Galileo Open Service Navigation Message Authentication (OSNMA) officially became operational in July of 2025 and utilizes a combination of the TESLA protocol and digital signature to authenticate satellite positioning messages (european_spce_agency_galileo_2021; european_gnss_supervisory_authority_galileo_2022). Furthermore, the TESLA authentication protocol has been proposed as a means to authenticate various types of broadcast services, including Automated Identification System (AIS) for marine traffic information broadcast (sciancalepore_auth-ais_2022), as well as Automated Dependent Surveillance Broadcast (ADS-B) systems for air traffic information broadcast (berthier_sat_2017). More recently, it was identified as a potential solution for providing authentication for the very-high-frequency data exchange system (VDES) that is anticipated to enhance a wide variety of maritime communications capabilities  (wimpenny_pragmatic_2025). We are the first to propose using TESLA for Remote ID broadcast authentication.
@@
TEE capabilities: TEEs in mobile computing are secure, isolated environments where the most critical operations related to key generation and storage can be safely executed (jauernig_trusted_2020). Many mobile applications are becoming increasingly reliant on the services provided by TEEs. They are used to provide services ranging from Mobile Network Operator (MNO) billing, digital rights management, and biometric authentication (paju_sok_2023; ahmad_enhancing_2013). In our specific application, the TEE’s integrity and secure storage capabilities are leveraged to conduct the transmission setup procedures required by the TESLA protocol.
3. Design of TBRD
@@
3.1. Threat Model and Assumptions
Our threat model addresses adversaries who have radio access to Remote ID broadcasts and aim to exploit the absence of authentication to inject false or misleading Remote ID broadcasts. We assume an attacker can either passively observe and/or actively transmit wireless data with basic SDR capabilities (e.g., USRP or HackRF). Furthermore, they have sufficient knowledge to record and replay Wi-Fi beacons or Bluetooth advertisements, the two FAA-approved broadcast media. We also consider attackers who can craft and transmit arbitrary Remote ID messages using open-source toolchains, including those that comply with the Remote ID standard but contain spoofed data. For example, an attacker could capture Remote ID broadcasts from a legitimate UAS mission and replay them at a later time to create the appearance of unauthorized activity. During FAA-designated spectator or sporting events, temporary flight restrictions may be in place over stadiums and other venues that normally fall in open airspace. The attacker replays previously captured Remote ID messages in the restricted zone, making it appear as though a registered operator is conducting an illegal flight. This can trigger enforcement actions or initiate investigations against legitimate operators. In coordinated attacks, repeated replays across multiple zones could be used to overwhelm response efforts and misdirect law enforcement attention.
Figure 1. An attacker intercepts Remote ID broadcasts from a legitimate UAS to include Basic ID, Location/Vector, and Operator ID messages. The attacker uses these legitimate transmissions to mask its digital identity in an effort to confuse and disrupt counter-UAS operations.: Spoofed Remote ID Attack
@@ 3.1. Threat Model and Assumptions
Consider a scenario depicted in figure 1 where an adversary wishes to operate a contraband-filled drone into a restricted airspace (e.g., prison) monitored by an observer, such as a law enforcement Counter UAS (C-UAS) system. The adversary’s drone simply broadcasts standard-compliant Remote ID messages using either a fabricated or cloned identifier. Because the current standard leaves authentication mechanisms unspecified, the observing C-UAS system has no way to determine whether a message truly originated from a legitimate UAS without implementing more complex detection mechanisms. This critical vulnerability enables simple spoofing, relay, and replay attacks, which can undermine surveillance efforts by allowing the adversary’s drone to appear compliant, potentially delaying or preventing an intercept.
In contrast, our proposed TBRD system is designed to thwart such spoofing attacks by default. Furthermore, as we will demonstrate, TBRD’s architecture provides specific countermeasures even against more advanced threats, such as capture of a UAS and extraction of mission secret keys from insecure flash storage.
3.2. System Overview